# Synthesis and stability of biomolecules under Earth’s upper mantle conditions

### v1

@@ Abstract
How life started on Earth is a long-time unsolved mystery. There are various hypotheses ranging from outer space to seabed. Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions of NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e. 10 - 13 GPa, 1000 - 1400 K). Contrary to the previous assumption that larger organic molecules might readily dissociate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, and uracil-like molecules. Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Moreover, our findings support the “RNA world” hypothesis, as we observed the exclusive formation of the 5-membered-ring form of ribose. By exploring the depths of Earth’s interior, we have uncovered a previously unexplored pathway through which life may have originated. These findings have contributed to our evolving understanding of the fundamental conditions necessary for life to arise on our planet.
@@ I Introduction
The origin of life on Earth remains a profound scientific mystery. Numerous hypotheses have been put forward, but none have been conclusively verified perez2020prebiotic . As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential ”primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, leads to the abiotic origin of life on early Earth. In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production . Inspired by the ”primordial soup” theory, various prebiotic conditions have been proposed for the origin of life, ranging from the interstellar medium (ISM) herbst1995chemistry or other planets burton2012understanding ; sleep1998refugia to deep-sea hydrothermal vents takai2006ultramafics .
@@
Hydrothermal vents are fissures in Earth’s oceanic crust where hot water rich in gases and minerals is released. Those deep sea vents are considered to have suitable conditions for origin of life. The pressure (P) of hydrothermal vents is tens of MPa depending on the depth of water, and the temperature (T) can reach 700 K. It is widely believed that the origin of life is unlikely to occur in the deeper regions of the Earth, such as the upper mantle, where the pressure can reach up to $\sim$13 GPa, accompanied by temperatures of $\sim$1700 K; the extremely high P-T conditions could readily degrade crucial biomolecules in aqueous solutions miller1988submarine . Otake et al. studied the stability of amino acids and their oligomerization at 10–5.0 GPa and 453–673 K, and found that high pressure inhibited the decomposition of amino acids ; however, their experiments were conducted at the water-poor condition otake2011stability . Previously, the fluids in Earth’s mantle were commonly modelled as simple mixtures of small volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model . However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; sverjensky2020changing ; pan2016fate ; stolte2019large ; stolte2021water . These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules.
@@
It is very challenging to study the origin of life in laboratory settings due to the extreme environments involved. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry within Earth’s upper mantle remains an uncharted territory.
@@
Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found at the bottom of Earth’s upper mantle. We discovered a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including glycine, ribose, and uracil-like molecules. Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Our study has provided valuable insights into a new potential avenue for the origin of life within Earth’s interior, expanding our understanding of the conditions necessary for life to emerge.
@@ II Results and Discussion
We initially mixed the H2O, H2, CO, and NH3 molecules in the simulation box. The molecular number of each molecule is 15. We conducted AIMD simulations at three P-T conditions: 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K to see how chemical products vary with P and T. Those P-T conditions are typically found at the bottom of Earth’s upper mantle. After 400 ps simulations at each P-T condition, we found about 100 different organic species. Here we mainly focus on the formation of bio-relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1. Fig. 1 shows the chemical structures of the representative species. Species with more than five carbon atoms are excluded due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds. The number and percentage of these CN-containing species as functions of time under three different P-T conditions are shown in Figures S5 and S6, respectively, which suggest that the concentrations of these species gradually stabilize over time.
@@
At 10 GPa and 1000 K, most of the organic products belong to three main categories, as shown in Fig. 2(a). The “C2N2” species account for more than half of the CN-containing species, followed by the “CN” species, and finally “C2N”. Formamide is the primary product in the “CN” species saladino2012formamide . It contains the peptide bond, which is a significant precursor for the production of more complex biomolecules, such as amino acids pietrucci2015formamide , nucleic acids, proteins, and even sugars saladino2012genetics ; halfen2011formation . The “C2N” species has the same CN-backbone (“-N-C-C-”) as that of glycine, suggesting that the “C2N” species may be a potential reactant for the formation of amino acids. At this P-T condition, some of the “C2N” species are not stable and can be transformed into the “C2N2” species by reacting with ammonia or ammonium ions, so the CN-backbone of “C2N2” is “-N-C-C-N-”, and one interesting molecule was identified as alpha-hydroxy-glycineamide ($\alpha$-HGA).
@@
With increasing the temperature to 1400 K along the isobar at 10 GPa, all the six categories of organic products at this P-T condition are observed. The “CN2” and “C4Nx” fractions are very small, while the other four species have similar fractions. Compared with 10 GPa and 1000 K, the portion of “C2N2” becomes smaller, while there are more “C2N” species. The “C2N” species formed here also share the same CN backbone (“-N-C-C-”) with glycine. Notably, in the “C2N” species, it is possible to form $\alpha$-hydroxyglycine, which can subsequently evolve into glycine through simple chemical reactions saitta2014miller . The “C2N2” species have the CN-backbone of“-N-C-C-N-”, including $\alpha$-HGA and oxamide. They are important organic molecules in the formation of the building blocks of peptides ligterink2019formation . Among the “CN” species, formamide remains the most abundant one. Additionally, the “CN2” species with the “-N-C-N-” backbone was observed, which did not typically form at 10 GPa and 1000 K. One representative product of this species is methanediamine (CH6N2) carballeira2001role , which was reported as a vital intermediate or precursor in the abiotic formation of nucleobases marks2022preparation .
@@
With increasing the pressure to 13 GPa at 1400 K, the “CN” species becomes dominating, and the “C2N2” species comes next. For the “CN” species, in addition to formamide, some other crucial precursors for producing more complex biomolecules are identified, such as carbamic acid tossell2010happens , isocyanic acid belson1982preparation , and formimidic acid pranata1995computational . The “C2N2” species exhibit a distinct CN-backbone compared to the two systems at 10 GPa, forming as “-C-N-C-N-”, which is present in the organic matter of allophanic acid. This kind of acid plays a vital role in the synthesis of biomolecules (e.g. uracil), and would be used in subsequent simulations. Surprisingly, the fraction of “C2N” species sharply decreases to a small value, adopting a new CN-backbone structure of “-C-N-C-”, as observed in N-formylformamide sanz2009effects . This structural change makes it challenging for “C2N” species formed here to evolve into glycine-like molecules or larger amino acids. In the “CN2” species, we have discovered the presence of urea molecules. Urea is closely related with biological processes, so it would also serve as a valuable reactant for investigating the reactions to larger biomolecules.
@@
For all the three P-T conditions, a common feature is that the “CN” and “C2N2” species are always the major compositions among the six CN-containing species. Fig. 2 (b) shows the lifetime of these CN-containing species. The lifetime distribution in Fig. 2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all the three P-T conditions, suggesting that they are relatively more stable than other species.
@@
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high . Thus, we first investigated the formation of formamide. Fig. 3 (a) provides an illustration of the distinct formation pathways of formamide at three P-T conditions. The complete formation processes are shown in Figure S7. At 10 GPa and 1000 K, the formation process began with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. When the H+ ion encountered a CO molecule, a reaction occurred, resulting in the formation of a C-H bond and an aldehyde group (-CHO). Subsequently, the aldehyde group reacted with the NH3 molecule, leading to the formation of a C-N bond, and ultimately, one N-H bond broke to produce formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first broke, leading to the formation of a dissociative amino group (-NH2). The amino group then reacted directly with a CO molecule, forming the C-N bond. Later, the water molecule dissociated again, providing a hydrogen atom that ultimately contributed to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond was initially formed through the reaction of NH3 and CO molecules. After a very short time, the dissociated hydrogen atom from the water molecule also bonded to the carbon atoms, forming a C-H bond. Finally, by releasing a hydrogen atom from the NH3 group, the formamide molecule was generated. Although the formation pathways differ at three P-T conditions, we observed two common features: first, the reactants always include CO, H2O, and NH3, and second, the reaction pathways must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
For the generation of biomolecules in living matter, the formation of C-N bonds is the key step. We further applied the advanced sampling method (see details in method section) to calculate the free energy profile of the C-N bond formation at three P-T conditions, as shown in Fig. 3(b). The collective variable is the distance between the carbon and nitrogen atoms. At 10 GPa, 1000 K and 13 GPa, 1400 K, the C-N bonds, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states. The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more favorable thermodynamically. The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively. As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition. The free energy result agrees with the data in Fig 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions. Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively. The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable. The C-N distances are larger than that of a typical C-N covalent bond, which strongly implies the presence of an additional atom connecting the two. This is consistent with a wider variety of organic species at these two P-T conditions. Our free energy calculations suggest that the C-N bond formation is favorable at 10 GPa and 1400 K. With increasing the pressure to 13 GPa and keeping the temperature at 1400 K, the carbon and nitrogen atoms tend to be bonded together by a third atom.
@@
After studying the basic chemical products at the three P-T conditions, we now turned to investigate the formation of large biomolecules as potential building blocks of life on Earth. Previous simulations have extensively studied glycine, the smallest amino acid, under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering . In our simulations at 10 GPa and 1400 K, we discovered one molecule very similar to glycine, called $\alpha$-hydroxyglycine. Fig. 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide. The chemical evolution suggests that, a hydrogen atom attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical. Subsequently, a CO2 molecule reacted with the protonated formamide, resulting in the production of a molecule with the “-N-C-C-” CN-backbone through forming a C-C bond. Finally, another hydrogen atom was captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule. In comparison to glycine, $\alpha$-hydroxyglycine differs by having one H atom and an “-OH” group instead of two H atoms connecting to the $\alpha$-C. Glycine can be obtained if the “-OH” group is replaced by a single H atom. During this process, the formation of the C-H bond is a crucial step. To identify appropriate P-T conditions for further advancing the chemical evolution of glycine, we calculated the free energy as a function of the C-H distance in Fig. 4 (b). Considering that the CN-backbone of the “C2N” species in the 13 GPa - 1400 K is “-C-N-C-”, very different from the CN-backbone in the glycine molecule, we did not consider this system. From the energy curves, we can see that the energy minima occurs at 1.07 Å for both 10 GPa - 1000 K and 10 GPa - 1400 K, and the energy barriers for C-H breaking are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are quite similar. However, the change of free energy from C-H bond formation to C-H bond dissociation is more positive at 10 GPa - 1400 K (6.76 kcal/mol) compared with that of 10 GPa - 1000 K (2.63 kcal/mol), which illustrates that 10 GPa - 1400 K condition is still the favorable one for forming C-H bond.
To test if we can also obtain glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules and perform AIMD simulations at 10 GPa and 1400 K, as shown in Fig. 4(c). Initially, two $\alpha$-hydroxyglycine molecules would connect to a larger molecule, but later transformed into two new molecules after about 28 ps, dehydroglycine and glyoxylic acid. We mainly focus on dehydroglycine. As the simulation went on, this molecule possessed the amino group when one free proton was bonded to its N atom. After a shorter time, the $\alpha$-C also captured one proton, resulting in the final molecular structure of glycine at about 38 ps. Additionally, our observations indicate that once glycine is formed at 10 GPa - 1400 K, it remains at a stable state.
@@
Both proteins and DNA are widely recognized as the important biomolecules for building blocks of life. However, the question of which one came first in the emergence of life remains a topic of debate. This is due to the fact that the replication of DNA relies on the catalytic activity of enzymes, which are a type of protein. Conversely, the production of proteins also needs the guidance of DNA, resulting in a chicken-and-egg dilemma. To solve this problem, a famous hypothesis known as “The RNA world” was proposed, which believed that RNA may occur earlier than DNA and protein gilbert1986origin ; nisbet1986origin . During this stage, the RNA not only have the ability to reproduce themselves and carry genetic information, but also could catalyze certain biological reactions in life activities like enzymemcclain1987model ; forster1990external . So the synthesis of RNA plays an important role in the early stage of the origins of life. Generally, RNA is composed of ribose sugar, four kinds of nucleobases, and a phosphate group. Considering that our simulation systems only contain C, H, O, and N elements, we then also explore whether the ribose and nucleobases could be synthesized under extreme conditions.

### v2

# Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany

@@ Abstract
How life started on Earth is an unsolved mystery. There are various hypotheses for the location ranging from outer space to the seafloor, subseafloor or potentially deeper. Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions between NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e. 10 - 13 GPa, 1000 - 1400 K). Contrary to the previous assumptions that larger organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules. Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we observed the predominant formation of the five-membered-ring form of ribose at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA. We have uncovered a previously unexplored pathway through which the building blocks of biomolecules may have originated in early Earth and other planets. Our findings contribute to an evolving understanding of the fundamental conditions necessary for life to arise.
@@ I Introduction
The origin of life on Earth remains a profound scientific mystery. Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins ; sutherland2017opinion . As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential “primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, to form the first organic compounds. This leads to the abiotic origin of life on early Earth. In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production . Inspired by the “primordial soup” theory, various prebiotic environments have been proposed for the origin of life, ranging from the outer space ziurys2006chemistry ; sandford2020prebiotic ; krasnokutski2022pathway to deep-sea hydrothermal vents daniel2006origins ; martin2008hydrothermal .
@@
In those hypotheses, prebiotic chemistry typically happens at non-ambient conditions. The pressure (P) of hydrothermal vents, resulting from the interaction between seafloor and seawater producing hot fluids, is tens of megapascals depending on the depth of water, and the temperature (T) can reach 700 K. Those deep sea vents were considered to have suitable conditions for the origin of life martin2008hydrothermal , but previous studies also suggested that the high temperatures in the vents could readily degrade crucial biomolecules in aqueous solutions miller1988submarine . Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 ; however, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule . It was commonly assumed that large biomolecules could not persist at both high hydrostatic pressures and high temperatures as found in Earth’s upper mantle, where pressures can reach up to $\sim$13 GPa, accompanied by high temperatures of $\sim$1700 K. In recent years, Otake et al. studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers; however, their experiments were conducted at the water-poor condition otake2011stability . In aqueous solutions, there seems a strong thermodynamic drive towards the formation of amino acids dimers above 1.5 GPa and temperature above 423 K robinson2021quantifying , in agreement with some experimental observations pedreira2019spontaneous . Previously, the fluids in Earth’s mantle were commonly modeled as simple mixtures of small unreactive volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model . However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; pan2013dielectric ; sverjensky2020changing ; pan2016fate ; abramson2017water ; stolte2019large ; stolte2021water . These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules in Earth’s or other planets’ subsurface or even deeper.
@@
It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . Note that the shock compression duration was less than 10 ps, so their simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions. AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
@@
Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle. We discovered that a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including formamide, glycine, ribose, urea, and uracil-like molecules. Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Our study provides valuable insights into a new potential avenue for the formation of building blocks of life and potentially for the origin of life within Earth’s and other planets’ interior, expanding our understanding of the conditions necessary for life to emerge.
@@ II Results and Discussion
The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Figure S1). Those molecules are typical volatile compounds found in C-H-O-N fluids in reducing environments of Earth’s deep interior zhang2009model ; manning2013chemistry , and are also abundant in interstellar space guelin2022organic . We conducted AIMD simulations at three P-T conditions, 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K, to evaluate how chemical products vary with P and T. Those P-T conditions are typically found at the bottom of Earth’s upper mantle. After more than 400 ps simulations at each P-T condition, we found about 100 different organic species. Here, we mainly focus on the formation of biotically relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1. Fig. 1 shows the chemical structures of the representative species. Species with more than five carbon atoms were observed; however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds. The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Figure S3, which suggest that the concentrations of these species gradually stabilize over time.
@@
At 10 GPa and 1000 K, most of the organic products belong to three main categories, as shown in Fig. 2(a). The “C2N2” species accounts for more than half of the CN-containing species, followed by the “CN” species, and finally “C2N” species. Formamide is the primary product in the “CN” species saladino2012formamide . It contains the peptide bond, which is a significant precursor for the production of more complex biomolecules, such as amino acids pietrucci2015formamide , nucleic acids, proteins, and even sugars saladino2012genetics ; halfen2011formation . The “C2N” species has the same CN-backbone (“-N-C-C-”) as glycine, suggesting that the “C2N” species may be a potential reactant for the formation of amino acids. At this P-T condition, some of the “C2N” species are not stable and can be transformed into the “C2N2” species by reacting with ammonia or ammonium ions, so the CN-backbone of “C2N2” species is “-N-C-C-N-”, and one interesting molecule is identified as alpha-hydroxy-glycineamide ($\alpha$-HGA).
@@
With increasing the temperature to 1400 K along the isobar at 10 GPa, all the six categories of organic products at this P-T condition were observed. Fractions of the “CN2” and “C4Nx” species are very small, while the other four species have similar fractions. Compared with 10 GPa and 1000 K, the portion of “C2N2” species becomes smaller, while there are more “C2N” species. The “C2N” species formed here also shares the same CN-backbone (“-N-C-C-”) with glycine. Notably, in the “C2N” species, we found a molecule of $\alpha$-hydroxyglycine, which can easily evolve into glycine saitta2014miller . The “C2N2” species has the CN-backbone of “-N-C-C-N-”, including $\alpha$-HGA and oxamide. They are important organic molecules in the formation of the building blocks of peptides ligterink2019formation . Among the “CN” species, formamide remains the most abundant one. Additionally, the “CN2” species with the “-N-C-N-” backbone was observed, which did not typically form at 10 GPa and 1000 K. One representative product of this species is methanediamine (CH6N2) carballeira2001role , which was reported as a vital intermediate or precursor in the abiotic formation of nucleobases marks2022preparation .
@@
With increasing the pressure to 13 GPa at 1400 K, the “CN” species becomes dominating, and the “C2N2” species comes next. For the “CN” species, besides formamide, some other crucial precursors for producing more complex biomolecules are identified, such as carbamic acid tossell2010happens , isocyanic acid belson1982preparation , and formimidic acid pranata1995computational . The “C2N2” species exhibits a distinct CN-backbone compared to the two solutions at 10 GPa, forming as “-C-N-C-N-”, which is present in allophanic acid. This acid plays a vital role in the synthesis of biomolecules, e.g., uracil, and would be used as reactants in subsequent simulations of forming larger biomolecules. Surprisingly, the fraction of “C2N” species sharply decreases, adopting a new CN-backbone structure of “-C-N-C-”, as observed in N-formylformamide sanz2009effects . This structural change makes it challenging for “C2N” species formed here to evolve into glycine-like molecules or larger amino acids. In the “CN2” species, we discovered the presence of urea molecule. Urea is closely related with biological processes daniel2006origins and considered to be an important condensation agent in prebiotic chemistry burcar2019stark ; gull2020silicate , so it would also serve as a valuable reactant for investigating the reactions to larger biomolecules.
@@
For all the three investigated P-T conditions, a common feature is that the “CN” and “C2N2” species are always the major components among the six CN-containing species. The lifetime distribution in Fig. 2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all three P-T conditions, suggesting that they are relatively more stable than other species.
@@
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high . Thus, we first investigated the formation of formamide molecule. Fig. 3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4. At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O). Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2). The amino group then reacts with a CO molecule, forming the C-N bond. Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules. After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond. Finally, by releasing a proton from the NH3 group, the formamide molecule is generated. Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
For the generation of biomolecules in living matter, the formation of C-N bonds is the key step. We further applied the advanced sampling method (see details in the Methods section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig. 3(b). The collective variable (CV) is the distance between the carbon and nitrogen atoms. Our calculations show that the C-N bonds at 10 GPa, 1000 K and 13 GPa, 1400 K, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states. The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more thermodynamically favored. The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively. As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition. The free energy result agrees with the data in Fig. 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions. Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively. The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable. The C-N distances are larger than that of a typical C-N covalent bond, which implies the presence of an additional atom connecting the two. This is consistent with a wider variety of organic species at these two P-T conditions. Overall, our free energy calculations suggest that the C-N bond formation is favored at 10 GPa and 1400 K.
@@
After studying the basic chemical products at three P-T conditions, we now turn to investigate the formation of large biomolecules as potential building blocks of life. Previous simulations have extensively studied glycine and the smallest amino acid under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering . In our simulations at 10 GPa and 1400 K, we discovered $\alpha$-hydroxyglycine, very similar to glycine. Fig. 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide. A proton attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical. Then, a CO2 molecule reacts with the protonated formamide, resulting in the “-N-C-C-” backbone through forming a C-C bond. Finally, another proton is captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule. In comparison to glycine, $\alpha$-hydroxyglycine differs by having one hydrogen atom and an “-OH” group instead of two H atoms connecting to $\alpha$-C. Glycine can be obtained if the “-OH” group is replaced by a hydrogen atom. During this process, the formation of the C-H bond is a crucial step. To explore appropriate P-T conditions for the possible synthesis of glycine, we calculated the free energy as a function of the C-H distance in Fig. 4 (b). Considering that the CN-backbone of the “C2N” species at 13 GPa - 1400 K is “-C-N-C-”, different from the CN-backbone in the glycine molecule, we did not consider this P-T condition. We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are very similar. After the C-H bond dissociation, the free energy increases more at 10 GPa and 1400 K (6.76 kcal/mol) than at 10 GPa and 1000 K (2.63 kcal/mol), indicating that the C-H bond is more favored at the former P-T condition.
To study the reaction from $\alpha$-hydroxyglycine to glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules, as shown in Fig. 4(c) and performed AIMD simulations at 10 GPa and 1400 K. Initially, two $\alpha$-hydroxyglycine molecules combine to form one molecule, but later dissociate to dehydroglycine and glyoxylic acid after $\sim$28 ps. As the simulation goes on, the C=N double bond in the dehydroglycine molecule receives two protons and finally dehydroglycine becomes glycine at $\sim$ 38 ps.
@@
Both proteins and DNA are widely recognized as the important biomolecules for building blocks of life. However, the question of which one came first in the emergence of life remains a topic of debate. This is due to the fact that the replication of DNA relies on the catalytic activity of enzymes, which are a type of proteins. Conversely, the production of proteins also needs the guidance of DNA, resulting in a chicken-and-egg dilemma. To solve this problem, a hypothesis known as “the RNA world” was proposed, which believed that RNA may occur earlier than DNA and protein gilbert1986origin ; nisbet1986origin . During this stage, the RNA not only have the ability to reproduce themselves and carry genetic information, but also could catalyze certain biological reactions in life activities like enzymes mcclain1987model ; forster1990external . Thus, the synthesis of RNA plays an critical role in the early stage of the origin of life. RNA is composed of ribose sugar, four kinds of nucleobases, and a phosphate group. Considering that our simulations only involve the carbon, hydrogen, oxygen, and nitrogen elements, we then explored whether the ribose and nucleobases could be synthesized under extreme conditions.